# On the Theoretical Foundation of Sparse Dictionary Learning in Mechanistic Interpretability

### v2

@@ Abstract
As AI models achieve remarkable capabilities across diverse domains, understanding what representations they learn and how they process information has become increasingly important for both scientific progress and trustworthy deployment. Recent works in mechanistic interpretability have shown that neural networks represent meaningful concepts as directions in their representation spaces and often encode diverse concepts in superposition. Various sparse dictionary learning (SDL) methods, including sparse autoencoders, transcoders, and crosscoders, address this by training auxiliary models with sparsity constraints to disentangle these superposed concepts into monosemantic features. These methods have demonstrated remarkable empirical success but have limited theoretical understanding. Existing theoretical work is limited to sparse autoencoders with tied-weight constraints, leaving the broader family of SDL methods without formal grounding. In this work, we develop the first unified theoretical framework considering SDL as one optimization problem. We demonstrate how diverse methods instantiate the theoretical framework and provide rigorous analysis of the optimization landscape. We provide novel theoretical explanations for empirically observed phenomena, including feature absorption and dead neurons. We design the Linear Representation Bench, a benchmark that strictly follows the Linear Representation Hypothesis, to evaluate SDL methods with fully accessible ground-truth features. Motivated by our theory and findings, we develop feature achoring, a novel technique applicable for all SDL methods, to enhance their feature recovery capabilities.
@@ 1 Introduction
As artificial intelligence is increasingly deployed in high-stakes applications, from medical diagnosis to autonomous systems, the need for model interpretability has grown from an academic curiosity to a practical necessity (Lipton, 2017; Rudin, 2019). Numerous approaches have been proposed to peer inside the “black box,” including probing classifiers (Belinkov, 2021), attention visualization (Vaswani et al., 2023), concept bottleneck models (Koh et al., 2020), and feature attribution methods (Sundararajan et al., 2017). Among the efforts to enhance model interpretability, one line of research focuses on the structure of the model representations, pointing out that AI models often encode interpretable concepts as linear directions in superposition (Park et al., 2024; Elhage et al., 2022). As a consequence, individual neurons are typically polysemantic, responding to multiple unrelated concepts unlike their biological counterparts (Bricken et al., 2023; Templeton et al., 2024; Quiroga et al., 2005).
To address the polysemantic model representations, Sparse Dictionary Learning (SDL) has emerged as a promising paradigm (Sharkey et al., 2025). By training auxiliary models, such as sparse autoencoders (SAEs) (Cunningham et al., 2023), transcoders (Dunefsky et al., 2024), and crosscoders (Lindsey et al., 2024), that reconstruct neural activations through a sparsity-constrained bottleneck, SDL methods aim to identify the interpretable concepts underlying the polysemantic representations. These methods have achieved remarkable empirical success, scaling to frontier language models (Templeton et al., 2024; Gao et al., 2024) and enabling applications ranging from feature steering (Wang et al., 2025) to circuit analysis (O’Neill and Bui, 2024) and interpretable medical diagnosis (Abdulaal et al., 2024).
Despite these empirical advances, due to a lack of unified theoretical framework to analyze SDL methods, the field identifies various phenomena without rigorous theoretical explanation, including feature splitting and feature absorption (Chanin et al., 2025), the existence of dead neurons and the effectiveness of resampling (Bricken et al., 2023). Without theoretical grounding, the development of improved SDL methods remains largely empirical, potentially leaving highly effective techniques underexplored. While previous works on sparse coding and dictionary learning have established solid theoretical results (Spielman et al., 2012; Gribonval and Schnass, 2010), these results do not directly address SDL methods used in mechanistic interpretability. Recent work by Cui et al. (2025) has taken a key step by establishing necessary and sufficient conditions for identifiable sparse autoencoders under the linear representation hypothesis. However, their framework focuses exclusively on SAEs with tied-weight constraints, leaving other SDL methods without theoretical grounding.
@@
In this work, we develop a unified theoretical framework that formalizes SDL as a general optimization problem, encompassing various SDL methods (Bussmann et al., 2024, 2025; Tang et al., 2025b) as special cases. We demonstrate how these diverse methods instantiate our framework through different choices of input-output representation pairs, activation functions, and loss designs. We establish rigorous conditions under which SDL methods provably recover ground-truth interpretable features, characterizing the roles of feature sparsity, latent dimensionality, and activation functions. Through detailed analysis of the optimization landscape, we demonstrate that global minima correspond to correct feature recovery while spurious local minima exhibit un-disentangled polysemanticity, providing novel theoretical explanation for feature absorption (Chanin et al., 2025) and neuron resampling (Bricken et al., 2023). We design the Linear Representation Bench, a synthetic benchmark that strictly follows the Linear Representation Hypothesis, to evaluate SDL methods with fully accessible ground-truth features. Motivated by our theoretical insights, we propose feature anchoring that achieve improved feature recovery by constraining learned features to known anchor directions. Our contributions are as follows:
@@
  • We provide rigorous theoretical results on the optimization landscape and provide novel explanations for various phenomena observed empirically.
  • We design the Linear Representation Bench, a benchmark enabling controlled evaluation of SDL methods with fully accessible ground-truth features.
  • We propose feature anchoring that constrains learned features to known anchor directions, achieving improved feature recovery performance.
@@ 2.1 Input Distribution and Model Representation
Let $\mathcal{D}$ denote the distribution over possible inputs to a neural network. For example, $\mathcal{D}$ could be the distribution of natural images or the distribution of text sequences.
@@ 2.2 The Linear Representation Hypothesis
A model representation $\mathbf{x}_{p}\in\mathbb{R}^{n_{p}}$ satisfies the Linear Representation Hypothesis (LRH) if there exists a feature function $\mathbf{x}:\mathcal{D}\to\mathbb{R}^{n}$ and a feature matrix $W_{p}\in\mathbb{R}^{n_{p}\times n}$ such that:
@@
    Non-negativity: $\mathbf{x}(s)\geq 0$ component-wise.
@@
    In typical superposition scenarios, $M>0$ but remains small (close to zero), characterizing a regime where many interpretable features are compressed into fewer dimensions with bounded mutual interference.
  3. 3.
    Extreme Sparsity: The sparsity level $S$ approaches 1:
    $S\to 1$
  4. 4.
where $W_{p}[:,i]$ denotes the $i$-th column of $W_{p}$.
@@ 2.3 General Optimization Framework for SDL
A SDL model maps an input representation $\mathbf{x}_{p}\in\mathbb{R}^{n_{p}}$ to a target representation $\mathbf{x}_{r}\in\mathbb{R}^{n_{r}}$ through a two-layer architecture:
@@
    An encoder layer that maps $\mathbf{x}_{p}$ to a latent space:
    $\mathbf{x}_{q}=\sigma(W_{e}\mathbf{x}_{p})$ (1)
@@
    A decoder layer that maps the latents to $\mathbf{x}_{r}$:
    $\hat{\mathbf{x}}_{r}=W_{d}\mathbf{x}_{q}$ (2)
where $\mathbf{x}_{q}\in\mathbb{R}^{n_{q}}$, $W_{e}\in\mathbb{R}^{n_{q}\times n_{p}}$, $W_{d}\in\mathbb{R}^{n_{r}\times n_{q}}$, and $\sigma:\mathbb{R}^{n_{q}}\to\mathbb{R}^{n_{q}}$ is a sparsity-inducing activation function.
The SDL objective minimizes expected reconstruction error:
@@ 2.4 Instantiations: Existing SDL Methods
We now demonstrate how existing SDL methods instantiate our framework through adopting different choices of input-target pairs $(\mathbf{x}_{p},\mathbf{x}_{r})$ and activation functions $\sigma$, and proposing variants on the loss function $\mathcal{L}_{\text{SDL}}$.
@@
Transcoders. Transcoders (Dunefsky et al., 2024; Paulo et al., 2025) capture interpretable features in layer-to-layer transformations. Unlike SAEs, transcoders approximate the inputoutput function of a target component, such as an an MLP, using a sparse bottleneck. In our proposed theoretical framework, transcoders set $\mathbf{x}_{p}=\text{NN}_{\ell}(s)$ and $\mathbf{x}_{r}=\text{NN}_{\ell+1}(s)$, where $\text{NN}_{\ell}$ denotes the target component’s input, and $\text{NN}_{\ell+1}$ denotes its output (Figure 2).
Figure 2: Transcoder: encoder $W_{e}$ maps $\text{NN}_{\ell}(s)$ to sparse latents $\mathbf{x}_{q}$, decoder $W_{d}$ predicts $\text{NN}_{\ell+1}(s)$.
@@
The variants described above demonstrate that diverse SDL methods can be unified under our general framework (Definition 2.5) through specific choices of activation functions $\sigma$ and loss modifications. This motivates our theoretical approach: rather than analyzing each variant separately, we establish a key property for all these activation functions.
Key Activation Function Property. Our analysis focuses on activation functions satisfying:
@@
This property holds for ReLU, JumpReLU, Top-$k$, Batch Top-$k$, and their compositions, the primary sparsity mechanisms in practice. Notably, all these activation functions can be expressed in the form $\sigma_{\text{Jump}}(z;c)$ for some threshold $c\geq 0$:
$\sigma_{\text{Jump}}(z;c)_{i}=z_{i}\cdot\mathbb{1}(z_{i}>c)$ (11)
@@ 3 Theoretical Results
Our theoretical analysis bridges this gap by characterizing the SDL optimization landscape under the Linear Representation Hypothesis. Section 3.1 derives the necessary and sufficient conditions for the global minima of the optimization problem. Section 3.2 demonstrates the existence of spurious local minima and how they relate with polysemanticity. Section 3.3 discusses about feature absorption and hierarchical concepts. We provide full proofs in Appendix.
3.1 Global Minimum
Here we analyze the global minima of our proposed framework. Theorem 3.1 shows that $\mathcal{L}_{\text{SDL}}$ is well-approximated by an alternative loss on feature reconstructions, $\tilde{\mathcal{L}}_{\text{SDL}}$.
@@ 3.1 Global Minimum
Under Assumption 2.3 and Assumption 2.4, define the approximate Loss, $\tilde{\mathcal{L}}_{\text{SDL}}$, as:
@@
Theorem 3.2 then demonstrates that the configuration $(W_{d}^{*},W_{e}^{*})=([W_{r},\mathbf{0}],[W_{p}^{\top};\mathbf{0}])$ achieves approximate loss bounded by $O((1-S)M^{2})$, establishing it as a near-optimal solution that recovers the ground-truth feature structure.
Theorem 3.2 (Approximate Loss at Global Minimum Configuration).
@@
Notice that the necessary and sufficient conditions for $\tilde{\mathcal{L}}_{\text{SDL}}=0$ require $w_{r}^{d}=W_{d}\sigma(W_{e}w_{p}^{d})$ for all $d\in[n]$. This system of $n$ vector equations in the encoder-decoder parameters is underdetermined, admitting multiple solutions beyond $(W_{d}^{*},W_{e}^{*})$. Crucially, some solutions achieve zero reconstruction loss without recovering any ground-truth features, as demonstrated in Figure 4. This underdetermined nature motivates feature anchoring, a novel technique that applies additional constraints to guide optimization.
@@
3.2 Polysemanticity and Local Minima
We now establish that SDL optimization can converge to spurious local minima with loss strictly greater than the global optimum. We also point out the connection between local minima and un-disentangled polysemanticity.
Example 3.3 (Spurious Local Minimum).
Consider $n=n_{p}=n_{q}=n_{r}=2$, $\sigma=\sigma_{\text{Top-1}}\cdot\sigma_{\text{ReLU}}$, and extreme sparsity ($S\to 1$). Let $W_{e}=\begin{bmatrix}(w_{e}^{1})^{\top}\\ (w_{e}^{2})^{\top}\end{bmatrix}$ and $W_{d}=[w_{d}^{1},w_{d}^{2}]$. The approximate loss decomposes as:
$\begin{split}\tilde{\mathcal{L}}_{\text{SDL}}(W_{d},W_{e})=M_{1}\left\|w_{r}^{1}-W_{d}\sigma(W_{e}w_{p}^{1})\right\|^{2}\\ +M_{2}\left\|w_{r}^{2}-W_{d}\sigma(W_{e}w_{p}^{2})\right\|^{2}\end{split}$ (17)
where $M_{1},M_{2}>0$ are constants.
Consider a configuration where:
$\langle w_{e}^{1},w_{p}^{1}\rangle>\langle w_{e}^{2},w_{p}^{1}\rangle>0,\quad\langle w_{e}^{1},w_{p}^{2}\rangle>\langle w_{e}^{2},w_{p}^{2}\rangle>0$ (18)

### v3

# A Unified Theory of Sparse Dictionary Learning in Mechanistic Interpretability: Piecewise Biconvexity and Spurious Minima

@@ Abstract
As AI models achieve remarkable capabilities across diverse domains, understanding what representations they learn and how they encode concepts has become increasingly important for both scientific progress and trustworthy deployment. Recent works in mechanistic interpretability have widely reported that neural networks represent meaningful concepts as linear directions in their representation spaces and often encode diverse concepts in superposition. Various sparse dictionary learning (SDL) methods, including sparse autoencoders, transcoders, and crosscoders, are utilized to address this by training auxiliary models with sparsity constraints to disentangle these superposed concepts into monosemantic features. These methods are the backbone of modern mechanistic interpretability, yet in practice they consistently produce polysemantic features, feature absorption, and dead neurons, with very limited theoretical understanding of why these phenomena occur. Existing theoretical work is limited to tied-weight sparse autoencoders, leaving the broader family of SDL methods without formal grounding. We develop the first unified theoretical framework that casts all major SDL variants as a single piecewise biconvex optimization problem, and characterize its global solution set, non-identifiability, and spurious optima. This analysis yields principled explanations for feature absorption and dead neurons. To expose these pathologies under full ground-truth access, we introduce the Linear Representation Bench. Guided by our theory, we propose feature anchoring, a novel technique that restores SDL identifiability, substantially improving feature recovery across synthetic benchmarks and real neural representations.
@@ 1 Introduction
As artificial intelligence systems scale to frontier capabilities, understanding their internal mechanisms has become essential for safe deployment (Lipton, 2017; Rudin, 2019). A central insight from mechanistic interpretability is that neural networks encode interpretable concepts as linear directions in superposition (Park et al., 2024; Elhage et al., 2022), where individual neurons respond to multiple unrelated concepts—a phenomenon known as polysemanticity (Bricken et al., 2023; Templeton et al., 2024). To disentangle these superposed representations, Sparse Dictionary Learning methods, including sparse autoencoders (SAEs) (Cunningham et al., 2023), transcoders (Dunefsky et al., 2024), and crosscoders (Lindsey et al., 2024), have emerged as the dominant paradigm, achieving remarkable empirical success on frontier language models (Templeton et al., 2024; Gao et al., 2024) and enabling applications from feature steering (Wang et al., 2025) to circuit analysis (Marks et al., 2025) and medical diagnosis (Abdulaal et al., 2024).
Despite this empirical success across diverse applications, SDL methods consistently exhibit persistent failure modes: learned features remain polysemantic (Chanin et al., 2025), ”dead neurons” fail to activate on any data samples (Bricken et al., 2023), and ”feature absorption” occurs where one neuron captures specific sub-concepts while another responds to the remaining related concepts (Chanin et al., 2025). Practitioners have developed techniques to address these issues, including neuron resampling (Bricken et al., 2023), auxiliary losses (Gao et al., 2024), and careful hyperparameter tuning, yet these fixes remain ad-hoc engineering solutions without principled justification. Critically, these phenomena persist even with careful training on clean data, suggesting they are not mere implementation artifacts but reflect fundamental structural properties of the SDL optimization problem itself.
We argue that the root cause of these failure modes is the non-identifiability of SDL methods: even under the idealized Linear Representation Hypothesis, SDL optimization admits multiple solutions achieving perfect reconstruction loss, some recovering no interpretable ground-truth features at all, necessitating comprehensive theoretical analysis on SDL methods. While classical dictionary learning theory provides identifiability guarantees under strict conditions (Spielman et al., 2012; Gribonval and Schnass, 2010), and recent work establishes necessary conditions for tied-weight SAEs (Cui et al., 2025), no unified theoretical framework explains why diverse SDL methods, SAEs, transcoders, crosscoders, and their variants, systematically fail in predictable ways. Without theoretical grounding, the development of improved SDL methods remains largely empirical, potentially leaving highly effective techniques underexplored.
@@
In this work, we develop a unified theoretical framework that formalizes SDL as a general optimization problem, encompassing various SDL methods (Bussmann et al., 2024, 2025; Tang et al., 2025b) as special cases. We demonstrate how these diverse methods instantiate our framework through different choices of input-output representation pairs, activation functions, and loss designs. We establish rigorous conditions under which SDL methods provably recover ground-truth interpretable features, characterizing the roles of feature sparsity, latent dimensionality, and activation functions. Through detailed analysis of the optimization landscape, we demonstrate that global minima correspond to correct feature recovery and provide necessary and sufficient conditions for achieving zero loss. We establish the prevalence of spurious partial minima exhibiting un-disentangled polysemanticity, providing novel theoretical explanations for feature absorption (Chanin et al., 2025) and the effectiveness of neuron resampling (Bricken et al., 2023). We design the Linear Representation Bench, a synthetic benchmark that strictly follows the Linear Representation Hypothesis, to evaluate SDL methods with fully accessible ground-truth features. Motivated by our theoretical insights, we propose feature anchoring, a technique applicable to all SDL methods which achieves improved feature recovery by constraining learned features to known anchor directions.
Our main contributions are as follows:
@@
  • We theoretically prove that SDL optimization is biconvex, bridging mechanistic interpretability methods with traditional biconvex optimization theory.
  • We characterize SDL optimization landscape and prove its non-identifiability, providing novel explanations for various phenomena observed empirically.
  • We design the Linear Representation Bench, a benchmark with fully accessible ground-truth features, enabling fully transparent evaluation of SDL methods.
  • We propose a novel technique, feature anchoring, that can achieve largely improved feature recovery performance applicable for all SDL methods. We validate the effectiveness of feature anchoring with extensive experiments across diverse SDL methods and settings.
@@ 2.1 Input Distribution and Model Representation
Let $\mathcal{D}$ denote the distribution over possible inputs, $\mathcal{X}$, to a neural network. For example, $\mathcal{D}$ could be the distribution of natural images or the distribution of text sequences (Notations in Appendix B).
@@ 2.2 The Linear Representation Hypothesis
A model representation $\mathbf{x}_{p}\in\mathbb{R}^{n_{p}}$ satisfies the Linear Representation Hypothesis (LRH) if there exists a feature function $\mathbf{x}:\mathcal{X}\to\mathbb{R}^{n}$ and a feature matrix $W_{p}\in\mathbb{R}^{n_{p}\times n}$ such that:
@@
    Non-negativity: $\mathbf{x}(s)\in\mathbb{R}^{n}_{+}$.
@@
where $W_{p}[:,i]$ denotes the $i$-th column of $W_{p}$. In typical superposition scenarios, $M>0$ but remains small (close to zero), characterizing a regime where many interpretable features are compressed into fewer dimensions with bounded mutual interference.
@@ 2.3 General Optimization Framework for SDL
A SDL model maps an input representation $\mathbf{x}_{p}(s)\in\mathbb{R}^{n_{p}}$ to a target representation $\mathbf{x}_{r}(s)\in\mathbb{R}^{n_{r}}$ through a two-layer architecture:
@@
    An encoder layer that maps $\mathbf{x}_{p}(s)$ to a latent space:
    $\mathbf{x}_{q}(s)=\sigma(W_{E}\mathbf{x}_{p}(s))$ (1)
@@
    A decoder layer that maps the latents to $\mathbf{x}_{r}(s)$:
    $\hat{\mathbf{x}}_{r}(s)=W_{D}\mathbf{x}_{q}(s)$ (2)
where $\mathbf{x}_{q}(s)\in\mathbb{R}^{n_{q}}$, $W_{E}\in\mathbb{R}^{n_{q}\times n_{p}}$, $W_{D}\in\mathbb{R}^{n_{r}\times n_{q}}$, and $\sigma:\mathbb{R}^{n_{q}}\to\mathbb{R}^{n_{q}}$ is a sparsity-inducing activation function.
The SDL objective minimizes mean square error:
@@ 2.4 Instantiations: Existing SDL Methods
We now demonstrate how existing SDL methods instantiate our framework through adopting different choices of input-target pairs $(\mathbf{x}_{p},\mathbf{x}_{r})$ and activation functions $\sigma$, and proposing variants on the loss function $\mathcal{L}_{\text{SDL}}$ (See Appendix A).
@@
Transcoders. Transcoders (Dunefsky et al., 2024; Paulo et al., 2025) capture interpretable features in layer-to-layer transformations. Unlike SAEs, transcoders approximate the input-output function of a target component, such as a MLP, using a sparse bottleneck. In our proposed theoretical framework, transcoders set $\mathbf{x}_{p}=\mathbf{x}_{\text{mid}}(s)$ and $\mathbf{x}_{r}=\mathbf{x}_{\text{pre}}(s)$, where $\mathbf{x}_{\text{mid}}(s)$ denotes the inputs of one MLP block, and $\mathbf{x}_{\text{pre}}(s)$ denotes the prediction of MLP’s outputs (Figure 2).
Figure 2: Transcoder: encoder $W_{E}$ maps $\mathbf{x}_{p}(s)$ to sparse latents $\mathbf{x}_{q}(s)$, decoder $W_{D}$ gives $\mathbf{x}_{r}(s)$ as a prediction of MLP’s output.
@@
The variants described above demonstrate that diverse SDL methods can be unified under our general framework (Definition 2.5) through specific choices of activation functions $\sigma$ and loss modifications. Critically, this unification enables subsequent theoretical analysis to apply to the entire family of SDL methods, rather than just a single architecture. We also state a key property for SDL activation functions:
@@
This property holds for ReLU, JumpReLU, Top-$k$, Batch Top-$k$, and their compositions, the primary sparsity mechanisms in practice. Notably, all these activation functions can be expressed in the form $\sigma_{\text{Jump}}(z;c)$ for some threshold $c\geq 0$: $\sigma_{\text{Jump}}(z;c)_{i}=z_{i}\cdot\mathbf{1}(z_{i}>c)$.
@@ 3 Theoretical Results
Our theoretical analysis bridges this gap by characterizing the SDL optimization landscape under the Linear Representation Hypothesis. Section 3.1 establishes a loss approximation enabling tractable analysis (Theorem 3.1). Section 3.2 proves SDL exhibits piecewise biconvex structure within activation pattern regions (Theorem 3.2). Section 3.3 characterizes the global minimum and shows the optimization is underdetermined (Theorem 3.3 and  3.4). Section 3.4 establishes the prevalence of spurious partial minima exhibiting polysemanticity (Theorem 3.7). Section 3.5 shows that hierarchical concept structures naturally induce feature absorption patterns that manifest as partial minima (Theorem 3.10). We provide full proofs in Appendix I.
3.1 Approximate Loss with Feature Reconstruction
Under extreme sparsity, the SDL loss decomposes into independent per-feature reconstruction terms.
@@ 3.1 Approximate Loss with Feature Reconstruction
Under Assumptions 2.3 and 2.4, define the approximate loss as:
@@
This approximation isolates the contribution of each ground-truth feature, making the optimization landscape amenable to convex analysis techniques applied subsequently.
3.2 SDL Optimization is Piecewise Biconvex
While the SDL loss is non-convex globally due to activation discontinuities, it exhibits favorable convex structure within regions of fixed activation patterns.
Theorem 3.2 (Bi-convex Structure of SDL).
Consider the approximate SDL loss $\tilde{\mathcal{L}}_{\text{SDL}}(W_{D},W_{E})$. Define the activation pattern region as $\Omega_{\mathcal{A}}=\{W_{E}\in\mathbb{R}^{n_{q}\times n_{p}}:\forall d\in[n],\mathcal{A}(d)=\{i\in[n_{q}]:\langle w_{E}^{i},w_{p}^{d}\rangle>c\}\}$ where $c\geq 0$ is the activation threshold and $\mathcal{A}(d)\subseteq[n_{q}]$ denotes neurons activated by feature $d$.
Then $\tilde{\mathcal{L}}_{\text{SDL}}$ exhibits bi-convex structure over $\mathbb{R}^{n_{r}\times n_{q}}\times\Omega_{\mathcal{A}}$:
  1. 1.
    For any fixed $W_{E}\in\Omega_{\mathcal{A}}$, $W_{D}\mapsto\tilde{\mathcal{L}}_{\text{SDL}}(W_{D},W_{E})$ is convex in $W_{D}$.
  2. 2.
    For any fixed $W_{D}\in R^{nr\texttimes nq},W_{E}\mapsto\tilde{\mathcal{L}}_{\text{SDL}}(W_{D},W_{E})$ is convex in $W_{E}$ over $\Omega_{\mathcal{A}}$.
This establishes SDL as a biconvex optimization problem, bridging mechanistic interpretability with classical biconvex optimization theory. This characterization enables both theoretical analysis of the optimization landscape and implementation of specialized biconvex methods for SDL.
3.3 Characterizing the Global Minimum of SDL
Characterizing the global minimum reveals both the success conditions for SDL training and the fundamental underdetermined nature of the optimization problem.
Theorem 3.3 (Successful Reconstruction Achieves Near-Zero Loss).
@@ 3.3 Characterizing the Global Minimum of SDL
Theorem 3.4 (Necessary and Sufficient Conditions for Zero Loss).
The approximate loss satisfies $\tilde{\mathcal{L}}_{\text{SDL}}(W_{D},W_{E})=0$ if and only if
$w_{r}^{d}=W_{D}\sigma(W_{E}w_{p}^{d})\quad\text{for all }d\in[n]$ (16)
Theorem 3.3 provides a constructive global minimum that recovers ground-truth features, while Theorem 3.4 reveals the complete solution space. This system of $n$ vector equations is underdetermined when $n_{q}>n$, admitting multiple solutions beyond the feature-recovering configuration. Critically, some solutions achieve zero reconstruction loss without recovering any ground-truth features (Figure 4).
@@
3.4 Characterizing Spurious Partial Minima of SDL
Beyond the global minimum, SDL optimization exhibits spurious partial minima where neurons exhibit polysemanticity—responding to multiple unrelated features.
Example 3.5 (Spurious Partial Minimum).
Consider $n=n_{p}=n_{q}=n_{r}=2$, $\sigma=\sigma_{\text{Top-1}}\cdot\sigma_{\text{ReLU}}$, $S\to 1$, and $M_{1}=M_{2}=1$. Let:
$w_{p}^{1}=\begin{bmatrix}1\\ 0\end{bmatrix},\quad w_{p}^{2}=\begin{bmatrix}0\\ 1\end{bmatrix},\quad w_{r}^{1}=\begin{bmatrix}1\\ 0\end{bmatrix},\quad w_{r}^{2}=\begin{bmatrix}0\\ 1\end{bmatrix}$ (17)